OpenStudy (anonymous):

Prove the identity tan^2 u(1+cos 2u)=1-cos 2u

OpenStudy (rogue):

Rewrite what with only sin x and cos x?

OpenStudy (anonymous):

I just put the rest up there

OpenStudy (anonymous):

2sinxcosx + sinx =sinx( 2cosx + 1)

OpenStudy (anonymous):

\[\cos ^{2}x -\sin ^{2}x +sinx\]

OpenStudy (rogue):

Beketso, you mixed up cos 2x with sin 2x =P. Good job Myko! :)

OpenStudy (anonymous):

1+2sin^2x+sin x im lost very

OpenStudy (anonymous):

oh.pardom me

OpenStudy (anonymous):

np

OpenStudy (anonymous):

\[1-2\sin ^{2}x+sinx\]

OpenStudy (anonymous):

Thank you so much

OpenStudy (rogue):

\[\cos 2x = \cos^2 x - \sin^2 x, \sin^2 x + \cos^2 x = 1\]\[\sin^2 x = 1 - \cos^2 x, \cos^2 x = 1 - \sin^2x\] \[\cos 2x + \sin x = \cos^2 x - \sin^2 x + \sin x \]\[\cos 2x + \sin x = (1 - \sin^2 x) - \sin^2 x + \sin x = 1 - 2\sin^2 x + \sin x\]

OpenStudy (anonymous):

Prove the identity cos(x)/sec(x)-1 - cos(x)/sec(x)+1=2 cos(x)/tan^2(x)

OpenStudy (anonymous):

Prove the identity tan^2 u(1+cos 2u)=1-cos 2u

OpenStudy (anonymous):

Prove the identity tan^2 u(1+cos 2u)=1-cos 2u